# Preference-based Centrality and Ranking in General Metric Spaces

### v1

@@ 4 Simulation Studies
This section investigates the finite-sample behavior of the proposed CORE estimators under Euclidean geometry. The simulations are designed to address two complementary questions. First, in a one-dimensional setting where the population preference structure is analytically tractable, we examine how the estimated score functions converge as the sample size increases and compare the gradient-based and spectral implementations. Second, in moderate and high dimensions, we evaluate the ability of CORE to recover the population benchmark ranking induced by the preference centrality $r_{F}(\cdot)$ and contrast its performance with classical depth-based procedures.
@@ 4.1 One-dimensional score convergence
We consider sample sizes $n\in\{50,200,500,2000\}$ with $R=20$ independent replications for each $n$. In each replication, we generate $x_{1},\ldots,x_{n}\overset{\mathrm{i.i.d.}}{\sim}F$, where $F$ is either (i) the standard normal distribution $N(0,1)$, or (ii) a skewed Laplace distribution with density
@@
We compare three constructions of the pairwise preference matrix. First, we consider the population matrix $\mathbf{P}=\{p_{ij}\}$ with $p_{ij}=p_{F}(x_{i},x_{j})$. In one dimension, letting $F_{\mathrm{cdf}}$ denote the distribution function of $F$, the preference probability admits the closed form
@@
Second, we use the independent reference-sample estimator $\mathbf{P}^{\mathrm{ref}}=\{\hat{p}_{ij}^{\mathrm{ref}}\}$ defined in (8), computed from an independent sample $Z_{1},\ldots,Z_{n}\overset{\mathrm{i.i.d.}}{\sim}F$. Third, we consider the leave-two-out estimator $\mathbf{P}^{\mathrm{lto}}=\{\hat{p}_{ij}^{\mathrm{lto}}\}$ in (7), which uses the observed sample itself as reference points.
For each preference matrix, we compute the corresponding score vector using the gradient-based implementation (Algorithm 1), referred to as Population-GD, Reference-GD, and Leave-out-GD, respectively. In addition, we apply the spectral estimator (Algorithm 2) to $\mathbf{P}^{\mathrm{lto}}$, yielding the Leave-out-Spectral estimator.
Figure 1: One-dimensional comparison of four CORE estimators. Each panel corresponds to a distribution (Normal or skewed Laplace) and a sample size $n\in\{50,200,500,2000\}$. Estimated scores $\hat{\theta}_{i}$ are plotted against the sample points $x_{i}$. The vertical dashed line marks the population maximizer $\mu^{\star}$ of $r(\cdot)$.
@@
Figure 3: Mean Pearson correlation between the estimated score vectors and the Population-GD benchmark over $R=20$ independent replications. Points indicate the mean correlation and vertical error bars denote $\pm 1$ standard deviation.
Figure 2 fixes the distribution and estimation method and overlays the score curves across different sample sizes. For all methods, the curves stabilize toward a common limiting shape as $n$ grows. The vertical dashed line marks the population maximizer $\mu^{\star}$ of $r(\cdot)$ ($\mu^{\star}=0$ for the normal distribution and $\mu^{\star}=-0.462$ for the skewed Laplace case). Even for moderate sample sizes, the estimated scores peak near $\mu^{\star}$, demonstrating reliable recovery of the population center.
@@ 4.2 Rank recovery in moderate and high dimensions
We next study the ability of CORE and competing procedures to recover the population ranking induced by the preference-based centrality functional $r_{F}(\cdot)$ under the Euclidean dissimilarity $\delta(x,y)=\|x-y\|_{2}$, in moderate to high dimensions. We consider four $(n,d)$ configurations: $(150,80)$ and $(500,200)$ representing the $n>d$ regime, and $(80,200)$ and $(150,500)$ representing the high-dimension, low-sample-size regime $n<d$. For each configuration, results are averaged over $R=20$ independent replicates.
Data are generated from three representative distributions: (i) a multivariate Student’s $t$ distribution with $\nu=5$ degrees of freedom and identity scatter; (ii) a balanced two-component Gaussian mixture $0.5\,N(u,I_{d})+0.5\,N(-u,I_{d})$, where $u=(4,\ldots,4,0,\ldots,0)^{\top}$ has value $4$ on the first $10$ coordinates; and (iii) a coordinate-wise skewed Laplace model $\mathrm{ALD}(0,2,10)$, with the first coordinate further scaled by a factor of $100$, introducing pronounced skewness and anisotropy.
In each replicate, we construct the leave-two-out pairwise probability matrix $\{\hat{p}_{ij}^{\mathrm{lto}}\}$ defined in (7). We then compute CORE-GD using Algorithm S.1 (Section S.2), and CORE-Spectral using Algorithm 2. For CORE-GD, we report the centered score vector $\widehat{\boldsymbol{\theta}}$, while for CORE-Spectral we report the stationary score vector $\widehat{\mathbf{s}}$.
@@
Table 1: Spearman correlation ($\text{Mean}_{\text{SD}}$ over $R=20$ replicates) between each method’s score and Monte Carlo $\hat{r}_{F}(X_{i})$ for $(n,d)\in\{(150,80),(500,200)\}$.
@@
Table 2: Spearman correlation ($\text{Mean}_{\text{SD}}$ over $R=20$ replicates) between each method’s score and Monte Carlo $\hat{r}_{F}(X_{i})$ for the $d>n$ settings ($(n,d)\in\{(80,200),(150,500)\}$).
@@
Table 3: Spearman correlation ($\text{Mean}_{\text{SD}}$ over $R=20$ replicates) between each method’s score and the true log-density $\log f(X_{i})$ for $(n,d)\in\{(150,80),(500,200)\}$.
@@
Table 4: Spearman correlation ($\text{Mean}_{\text{SD}}$ over $R=20$ replicates) between each method’s score and the true log-density $\log f(X_{i})$ for $(n,d)\in\{(80,200),(150,500)\}$.
@@ 5 Conclusion and Discussion
We proposed CORE, a preference-based framework for defining population centrality through reference-based pairwise comparisons on a general dissimilarity space $(\mathcal{X},\delta)$. By projecting the induced pairwise preference structure onto a one-dimensional BTL score, CORE aggregates local comparison information into a globally interpretable ranking. A central conceptual contribution of this work is the population functional $r_{F}(\cdot)$, which provides an intrinsic notion of preference-based centrality that is defined independently of any parametric scoring model. Because $r_{F}(\cdot)$ depends explicitly on the chosen dissimilarity $\delta$, the framework offers a flexible mechanism for tailoring centrality to the geometry and semantics of the problem at hand.
On the methodological side, we developed two complementary estimators. The optimization-based estimator, CORE-GD, solves the empirical M-projection via first-order constrained optimization and yields calibrated scores with strong statistical and numerical stability. To address scalability concerns, we also introduced CORE-Spectral, a fast spectral approximation based on a Markov chain construction, which retains the essential ranking structure while significantly reducing computational cost. Our theoretical analysis establishes existence, uniqueness, and intrinsic monotonicity of the population target, and shows that the induced score satisfies the canonical axioms of statistical depth under mild regularity conditions.

### v2

@@ 4 Simulation Studies
This section investigates the finite-sample behavior of the proposed CORE estimators under Euclidean geometry. The simulations are designed to validate the structural and statistical properties established in Section 3. In particular, we examine: (i) whether the empirical estimators recover the population M-projection $\theta^{\star}$, (ii) whether the induced ranking aligns with the intrinsic preference centrality $r_{F}(\cdot)$, and (iii) how CORE compares with classical depth-based procedures across different geometric regimes.
@@ 4.1 One-dimensional score convergence
We consider sample sizes $n\in\{50,200,500,2000\}$ with $20$ independent replications for each $n$. In each replication, we generate $x_{1},\ldots,x_{n}\overset{\mathrm{i.i.d.}}{\sim}F$, where $F$ is either (i) the standard normal distribution $N(0,1)$, or (ii) a skewed Laplace distribution with density
@@
We compare three constructions of the pairwise preference matrix. First, we consider the population matrix $\mathbf{P}^{\mathrm{pop}}=\{p_{ij}\}_{i,j=1}^{n}$ with $p_{ij}=p_{F}(x_{i},x_{j})$. In one dimension, letting $F_{\mathrm{cdf}}$ denote the distribution function of $F$, the preference probability admits the closed form
@@
Second, we use the independent reference-sample estimator $\widehat{\mathbf{P}}^{\mathrm{ref}}=\{\hat{p}_{ij}^{\mathrm{ref}}\}_{i,j=1}^{n}$ defined in (8), computed from an independent sample $Z_{1},\ldots,Z_{n}\overset{\mathrm{i.i.d.}}{\sim}F$. Third, we consider the leave-two-out estimator $\widehat{\mathbf{P}}^{\mathrm{lto}}=\{\hat{p}_{ij}^{\mathrm{lto}}\}_{i,j=1}^{n}$ in (7), which uses the observed sample itself as reference points.
For each preference matrix, we compute the corresponding score vector using the gradient-based implementation (Algorithm 1), referred to as Population-GD, Reference-GD, and Leave-out-GD, respectively. In addition, we apply the spectral estimator (Algorithm 2) to $\widehat{\mathbf{P}}^{\mathrm{lto}}$, yielding the Leave-out-Spectral estimator.
Figure 1: One-dimensional comparison of four CORE estimators. Each panel corresponds to a distribution (Normal or skewed Laplace) and a sample size $n\in\{50,200,500,2000\}$. Estimated scores $\hat{\theta}_{i}$ are plotted against the sample points $x_{i}$. The vertical dashed line marks the population maximizer $\mu^{\star}$ of $r_{F}(\cdot)$.
@@
Figure 3: Mean Pearson correlation between the estimated score vectors and the Population-GD benchmark over $20$ independent replications. Points indicate the mean correlation and vertical error bars denote $\pm 1$ standard deviation.
Figure 2 fixes the distribution and estimation method and overlays the score curves across different sample sizes. For all methods, the curves stabilize toward a common limiting shape as $n$ grows. The vertical dashed line marks the population maximizer $\mu^{\star}$ of $r_{F}(\cdot)$ ($\mu^{\star}=0$ for the normal distribution and $\mu^{\star}=-0.462$ for the skewed Laplace case). Even for moderate sample sizes, the estimated scores peak near $\mu^{\star}$, demonstrating reliable recovery of the population center.
@@ 4.2 Rank recovery in moderate and high dimensions
We next evaluate the ability of CORE and competing procedures to recover the population ranking induced by the preference-based centrality functional $r_{F}(\cdot)$ under Euclidean dissimilarity $\delta(x,y)=\|x-y\|_{2}$ in moderate and high dimensions.
We consider four $(n,d)$ settings: $(150,80)$ and $(500,200)$ for the $n>d$ regime, and $(80,200)$ and $(150,500)$ for the high-dimensional regime $n<d$. All results are averaged over $20$ independent replicates.
Data are generated from three representative models: (i) a multivariate Student’s $t_{\nu}$ distribution with $\nu=5$ and identity scatter; (ii) a balanced Gaussian mixture $0.5,N(u,I_{d})+0.5,N(-u,I_{d})$, where $u=(4,\ldots,4,0,\ldots,0)^{\top}$ has value $4$ on the first $10$ coordinates; and (iii) a coordinate-wise skewed Laplace model $\mathrm{ALD}(0,2,10)$, with the first coordinate scaled by $100$ to induce strong skewness and anisotropy.
For each replicate, we construct the leave-two-out pairwise probability matrix $\widehat{\mathbf{P}}^{\mathrm{lto}}$ defined in (7). CORE-GD reports the centered score vector $\widehat{\boldsymbol{\theta}}$, whereas CORE-Spectral reports the stationary score vector $\widehat{\mathbf{s}}$.
@@
Table 1: Spearman correlation ($\text{Mean}_{\text{SD}}$ over 2020 replicates) between each method’s score and Monte Carlo $\hat{r}_{F}(X_{i})$ for $(n,d)\in\{(150,80),(500,200)\}$.
@@
Table 2: Spearman correlation ($\text{Mean}_{\text{SD}}$ over 2020 replicates) between each method’s score and Monte Carlo $\hat{r}_{F}(X_{i})$ for the $d>n$ settings ($(n,d)\in\{(80,200),(150,500)\}$).
@@
Table 3: Spearman correlation ($\text{Mean}_{\text{SD}}$ over 2020 replicates) between each method’s score and the true log-density $\log f(X_{i})$ for $(n,d)\in\{(150,80),(500,200)\}$.
@@
Table 4: Spearman correlation ($\text{Mean}_{\text{SD}}$ over 2020 replicates) between each method’s score and the true log-density $\log f(X_{i})$ for $(n,d)\in\{(80,200),(150,500)\}$.
@@
We also compare computation time across methods. CORE-Spectral is consistently faster than CORE-GD, which is expected since CORE-Spectral reduces to a single stationary-distribution computation for a Markov transition matrix, whereas CORE-GD solves a smooth convex program by iterative optimization. Among competing depth-based procedures, CORE-Spectral is typically faster than halfspace, projection, and spatial depths, whose implementations involve repeated optimization or many random projections. The only faster baselines are Mahalanobis depth and Neg-$\ell_{2}$, as they require only moment calculations (mean/covariance) and simple Euclidean norms. Detailed timing results are reported in Supplementary Section S.3.
5 Real Data Analysis
5.1 Application to the Madelon data
We study a dataset from the UCI Machine Learning Repository. [^1] https://archive.ics.uci.edu/dataset/171/madelon The dateset is generated from $32$ clusters positioned on the vertices of a five-dimensional hypercube with $n=2000$ and $p=500$. Only $20$ features are informative, with the remaining $480$ features serving as non-informative “probe” variables. We compute the score vector $\hat{\boldsymbol{\theta}}=(\hat{\theta}_{1},\ldots,\hat{\theta}_{n})$ using both CORE-GD and CORE-Spectral.
Figure 4 visualizes the resulting scores on a two-dimensional embedding. We apply PCA to the original $500$-dimensional data and project each sample onto the first two principal components. Each point is then colored by its estimated score $\hat{\theta}_{i}$, ranging from blue (high) through yellow to orange (low). The left and right panels correspond to CORE-GD and CORE-Spectral, respectively. In each panel, we mark the $10$ samples with the largest $\hat{\theta}_{i}$ values using green circles and the $10$ samples with the smallest $\hat{\theta}_{i}$ values using red circles. The two panels are nearly identical. The highest-score points concentrate near the center of the PCA plane, while the lowest-score points tend to lie toward the periphery. This pattern is consistent with the definition of preference-based centrality: points in denser regions of the induced dissimilarity geometry tend to be closer to a typical draw from the distribution and therefore achieve higher values of $r_{F}(\cdot)$, and hence higher calibrated scores via the monotone link in Theorem 2.
Figure 4: PCA visualization of the Madelon data. Left: CORE-GD. Right: CORE-Spectral. Points are colored by the estimated centrality score $\hat{\theta}_{i}$ from blue (high) through yellow to orange (low). The $10$ largest-score points are highlighted by green circles and the $10$ smallest-score points are highlighted by red circles.
5.2 Application to the Medfly data
We apply the proposed CORE methodology to the medfly25 dataset, which records daily egg-laying counts for female Mediterranean fruit flies (medflies, Ceratitis capitata) from the study of Carey et al. (1998). Each observation is an egg-laying trajectory observed over the early life span of a fly, and the dataset has been widely used as a benchmark for functional outlier detection. The preprocessed version provided in the fdapace R package contains $789$ trajectories observed over the first $25$ days after hatch. To assess stability across repeated experiments, we randomly draw subsamples of size $n=200$ and repeat the analysis over $100$ replicates. This subsampling strategy allows us to compute mean and standard deviation of rank correlations across methods.
In the simulation section, we compared CORE with several multivariate depth notions. However, those depth methods are designed for finite-dimensional Euclidean data and cannot be directly applied to functional trajectories. Therefore, in this real-data experiment we compare CORE with the regularized halfspace depth (RHD) proposed by Yeon et al. (2025), which is specifically developed for functional data.
For each subsample, we compute three centrality scores: CORE-GD, CORE-Spectral, and RHD. Pairwise dissimilarities between trajectories are defined using the $L^{2}$ metric. For two trajectories $x_{i}(t)$ and $x_{j}(t)$ observed over $\mathcal{T}$, we set
$\delta(x_{i},x_{j})=\left(\int_{\mathcal{T}}\{x_{i}(t)-x_{j}(t)\}^{2}\,dt\right)^{1/2},$
which is approximated numerically using the discretized observations over the first 25 days.
We first compare CORE-GD and CORE-Spectral. Table 5 reports the Spearman correlation between the two scores over $100$ random subsamples. The correlation is extremely high ($0.996$ with negligible variability), showing that the spectral approximation yields rankings almost identical to those obtained from the convex M-estimation formulation. This confirms that CORE-Spectral provides an accurate approximation to CORE-GD.
Table 5: Spearman correlation ($\text{Mean}_{\text{SD}}$ over $100$ replicates) between CORE-GD and CORE-Spectral scores on the medfly25 dataset.
Comparison Spearman
CORE-GD vs CORE-Spectral $0.996_{0.001}$
Since the two CORE variants produce nearly identical rankings, we focus on CORE-GD in the remainder of this section. Figure 5 visualizes the deepest trajectory and detected outliers based on CORE-GD scores for one subsample. The top panel displays all egg-laying trajectories, with the most central trajectory highlighted in green. Outliers are identified using the standard boxplot rule with coefficient $1.5$: an observation is flagged if its score falls below the lower fence $Q_{1}-1.5\times\mathrm{IQR}$, where $\mathrm{IQR}=Q_{3}-Q_{1}$. Using this rule, three trajectories (IDs 181, 61, and 197) are detected as outliers and highlighted in varying shades of red, with darker shades indicating lower centrality. The bottom panels display each flagged trajectory together with the deepest trajectory to illustrate their deviations in functional shape.
Table 6: Spearman correlation ($\text{Mean}_{\text{SD}}$ over $R=100$ replicates) between CORE-GD score and RHD depth on the medfly25 dataset, across truncation levels $J$ and quantile levels $u\in\{0.4,0.6,0.8,0.95\}$.
$u=0.4$ $u=0.6$ $u=0.8$ $u=0.95$
$J=6$ $0.813_{0.029}$ $0.780_{0.040}$ $0.736_{0.053}$ $0.692_{0.058}$
$J=8$ $0.803_{0.036}$ $0.774_{0.045}$ $0.731_{0.056}$ $0.689_{0.061}$
$J=10$ $0.801_{0.035}$ $0.765_{0.050}$ $0.726_{0.058}$ $0.683_{0.063}$
$J=15$ $0.795_{0.036}$ $0.758_{0.051}$ $0.717_{0.060}$ $0.676_{0.065}$
Figure 5: Centrality and outlier visualization of the medfly dataset based on CORE-GD scores. Three trajectories (IDs 181, 61, and 197) are detected as outliers by the boxplot rule and highlighted in varying shades of red. The top panel displays all curves, with the most central (deepest) trajectory marked in green, while the three detected outliers are separately shown in the bottom panels together with the deepest trajectory.
Finally, we compare CORE-GD with RHD, which involves two tuning parameters: the truncation level $J$, which controls the number of retained functional components, and the quantile level $u$, which determines the degree of regularization in the depth construction. To provide a comprehensive comparison, we evaluate RHD over a grid of parameter choices with $J\in\{6,8,10,15\}$ and $u\in\{0.4,0.6,0.8,0.95\}$. Table 6 reports the Spearman correlation between CORE-GD scores and RHD depths. The correlations range from $0.676$ to $0.813$, indicating substantial agreement in centrality rankings and supporting the validity of the proposed CORE framework in functional settings. Higher correlations are observed for smaller values of $J$ and $u$. Under these settings, RHD emphasizes global magnitude and coarse-scale structure, whereas larger $J$ and $u$ increase sensitivity to local shape features. Since CORE-GD aggregates global pairwise comparisons under the $L^{2}$ dissimilarity, its rankings align more closely with the globally oriented RHD configurations. Taken together, these findings suggest that CORE provides a complementary perspective to functional depth methods, achieving comparable centrality rankings while arising from a fundamentally different pairwise-comparison formulation.
@@ 6 Conclusion and Discussion
We proposed CORE, a preference-based framework for defining population centrality through reference-based pairwise comparisons on a general dissimilarity space $(\mathcal{X},\delta)$. Methodologically, we developed two complementary estimators, CORE-GD and CORE-Spectral, which together provide a statistically principled and computationally scalable approach to estimating preference-based rankings. Our theoretical analysis establishes existence and uniqueness of the population target, consistency of the proposed estimators, and shows that the induced ranking satisfies the canonical properties expected of a statistical depth notion.